OpenStudy (anonymous):

Organize the following expressions from greatest to least by number of terms: x + 2xyz 9x2yz 18x squared +5ab-6y 4x3+3x2-x-4 will give medal!!!

OpenStudy (anonymous):

The degree of a polynomial is the number of the highest exponent.

OpenStudy (anonymous):

9x2yz Degree=4 x=2 y=1 z=1 2+1+1=4 ,, x + 2xyz x x=1 -- 2xyz x=1 y=1 z=1 1+1+1=3 Degree=3 ,,, 2x3y + y2x – 3x + 4 2x^3y x=3 y=1 3+1=4 -- y^2x y=2 x=1 2+1=3 -- 3x x=1

OpenStudy (anonymous):

are you sure?